OpenStudy (lucky_luck):

find the volume of a cone that has a diameter of 14 cm. and a height of 9 cm. my answer came out to 461. am i correct? will medal and fan

OpenStudy (mathstudent55):

Correct, but be careful rounding off.

OpenStudy (mathstudent55):

\(\Large V_{cone} = \dfrac{1}{3} \pi r^2 h\) \( \Large = \dfrac{1}{3} \times \pi \times (7~cm)^2 \times (9~cm)\) \(\Large = 461.8141... ~cm^3\) Now round it off correctly.

OpenStudy (lucky_luck):

no cause the paper says round off to the nearest hundreth